OpenStudy (thegrimreaper666):

Would anyone care to refresh me on how to rationalize √2/9? I just need the steps, thanks.

OpenStudy (legomyego180):

Is this \[\sqrt{\frac{ 2 }{ 9 }}\]?

OpenStudy (thegrimreaper666):

Yep

OpenStudy (legomyego180):

\[\sqrt{\frac{ 2 }{ 9 }}=\frac{ \sqrt{2} }{ \sqrt{9} }=\frac{ \sqrt{2} }{ 3 }\]

OpenStudy (thegrimreaper666):

Right, thanks!